OpenStudy (anonymous):

Set both equations to y=mx+b x^2+y^2=17 x^2+y^2-2x=13 (sqrt-x^2 + 17 ) (-sqrt-x^2+17) (sqrt-x^2+2x+13) (-sqrt-x^2+2x+13) Is this right?

OpenStudy (amistre64):

circles arent normally linear...

OpenStudy (amistre64):

those may be more ovalish if anything, but same principle :)

OpenStudy (anonymous):

So its right though?

OpenStudy (amistre64):

cant tell, im still confused by the question :)

OpenStudy (anonymous):

im supposed to put the equations in mx+b so I can put it in my graphing calculator. Then im supposed to use the intersect function on my calculator to find the points. there should be 4 points.

OpenStudy (amistre64):

there are up to 4 possible points. that is right... but: y=mx+b is the equation for a line. What you have here are 2 ovals that might be intersecting adn there is no way to make any part of them a straight line that I am aware of

OpenStudy (amistre64):

your best bet would be to "eliminate" a variable first

OpenStudy (amistre64):

(x^2+y^2-0x=17) (-1) x^2+y^2-2x =13 -x^2 -y^2 +0x =-17 x^2 +y^2 -2x = 13 -------------------- 0x^2 +0y^2 -2x = -4 -2x = -4 x = 2 plug in x=2 for your x values and solve for y :)

OpenStudy (amistre64):

4 + y^2 = 17 y^2 = 13 y = sqrt(13) or y=-sqrt(13) 4 +y^2 -4 = 13 y^2 = 13 y= sqrt(13) or y=-sqrt(13) your points of intersection are: (2,sqrt(13)) and (2,-sqrt(13))

OpenStudy (anonymous):

Thank you!